OpenStudy (anonymous):

What is the LCD of (5)/(7x^3y^7) and (1)/(4x^4y^5)

OpenStudy (anonymous):

7x^3y^7 = 7*x*x*x*y*y*y*y*y*y*y 4x^4y^5 = 2*2*x*x*x*x*y*y*y*y*y

OpenStudy (anonymous):

Now, LCM = 7*4*x^4*y^7

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Alright..

OpenStudy (anonymous):

LCD of (5)/(7x^3y^7) and (1)/(4x^4y^5) = LCM of 7x^3y^7 and 4x^4y^5

OpenStudy (anonymous):

Alright.